What is the freezing temperature of an acetic acid/KCl mixture of concentration 0.55m if the Cryoscopic constant for acetic acid is 3.90 °C ·kg/mol and the normal freezing point of acetic acid is 16.6 oC? 


1
Expert's answer
2021-02-28T06:20:41-0500

"\Delta" Tf = Kf × mB

"\Delta T_f" = 3.90 × 0.55 = 2.14⁰C

(16.6 - T) = 2.14

T = 14.46⁰C